# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

### v2

@@ 6.1. Content Removal and Moderator Workload
Figure 5. For most topics, communities where moderators remove more content exhibit more negative sentiment (a). News communities, however, buck this trend, with the fraction of mod discourse with negative sentiment 11pp lower in news communities whose mods remove less than $1\%$ of content compared to communities whose mods remove 2% - 3% of content (b).
@@ 7.1. Limitations
Our work measures community governance through community members’ public discussion of moderators. While these signals enable insights about different governance strategies, they do not capture every aspect of the success of a community’s governance. What people say publicly does not always reflect their actual beliefs, and even if it did, minimizing community unhappiness is not necessarily the best objective function for community moderators. Differences in community members’ behavior and values also may bias our results, as different community may have different norms around mod discourse, and different community members may feel more or less comfortable expressing their opinions publicly. Additional study is needed to ensure that scalable methods for measuring governance reflect the needs of all community members, not just the noisiest.
@@ 7.2. Conclusion
Good governance is critical to the functioning of online communities, yet it is difficult to know what governance practices are most effective, as it is challenging to measure the ‘success’ of community governance. In this work, we developed a method to quantify community members’ perceptions of their moderators across thousands of communities. We relate these perceptions to different aspects of governance including community size and topic (§4) as well as to different actions that moderators can take, including rule enforcement (§6.1), community engagement (§6.2), and moderator recruiting practices (§6.3-6.4). We empirically identify promising strategies for community moderators, including tailoring the strictness of rule enforcement to the community topic, and recruiting engaged mods. We make our models and datasets public to support future research.
@@ 8. Ethical Considerations
We believe this work will have a positive broader impact by informing better moderation practices in online communities, as well as providing researchers with better tools to study community members’ perceptions of moderators. As we only make use of public data, we believe our work has minimal risk to participants’ privacy. As research has shown that some online community users are uncomfortable with their data being used for research, even when posted publicly (Fiesler and Proferes, 2018), we take further steps to reduce potential harms and misuse potential of our mod discourse dataset: we do not publish usernames or identifiable information, only predicted sentiment with regards to the moderators. We publish moderator timelines, including moderators’ usernames, however these usernames are already publicly listed on communities’ ‘about’ pages. Upon publication of this dataset, we will provide affordances for users to have their data removed at their request. This study was approved by the University of Washington IRB under ID number STUDY00011457.
Supplementary Materials
@@ A.3. Neutral Sentiment
Appendix B Supplementary Figures
@@
Appendix C Covariates Used in Propensity Score Modeling for IPTW
C.1. Covariates used for Mod Workload Analyses
@@
C.2. Covariates used for Strictness of Rule Enforcement Analyses

### v3

@@ 6.1. Content Removal and Moderator Workload
Figure 5. For most topics, communities where moderators remove more content exhibit more negative sentiment (a). News communities, however, buck this trend, with the fraction of mod discourse with negative sentiment 11 percentage poitns lower in news communities whose mods remove less than $1\%$ of content compared to communities whose mods remove 2% - 3% of content (b).
@@ 7. Discussion
Diversity of Communities’ Mod Perceptions. Our results show that communities’ perceptions of their moderators are highly varied (Figure 2 & 3) but are associated with other aspects of community health and size. This suggests that not only do different kinds of communities have different norms for their mod discourse, but also that community members’ perceptions of overall health of their community, which includes many factors outside of the moderators’ direct control, also influences their mod discourse. Researchers and moderators must carefully consider the specific needs of their communities.
Topic Contention Affects Governance. Our results suggest that the topic of a community impacts the community members’ perception of moderators’ removal of content and rule enforcement (Figure 5, Appendix Figure 10). We find that for news communities, as moderators remove more content, community members’ perception of moderators becomes more positive (up until $\approx 3\%$ of content is removed), while for other communities, perceptions of moderators are more negative as the mods remove more content. This suggests that communities with certain topics, perhaps more contentious ones, are more appropriate to moderate fairly strictly than others. Future work should specifically examine the impact of community topic on community members’ perceptions of content removal.
Community Overestimates of Moderators’ Power. There are many aspects of every community that moderators have limited-to-no control over. For example, our results show that perceptions of moderators are most positive in smaller communities (Figure 2f). However, although they can set and enforce policies with the goal of growing or shrinking a community, moderators cannot directly control a community’s size. Thus, our results support previous research that moderators are frequently ‘blamed’ for problems that they cannot easily resolve (Matias, 2019a). Future work could address this tension directly, perhaps by attempting to educate community members on the roles and powers of moderators, as well as the powers that moderators do not have.
Mod Recruiting and Engagement. We examine instances of moderators being added to mod teams to compute the impact that different types of moderators have on their communities. We find that more engaged moderators are associated with an increase in the amount of positive mod discourse, perhaps as a result of an increased sense of transparency and/or accountability for community governance (Figure 6a-c). We also look at differences between moderators whose recruiting was discussed publicly, versus those who were recruited in private. We find that publicly recruited moderators appear to be more controversial, with both negative and positive discourse increasing after their appointment (Figure 7b). This suggests that moderators must use caution when recruiting moderators publicly, and be sure to consider community preferences.
@@ 7.1. Limitations
Our work measures community governance through community members’ public discussion of moderators. While these signals enable insights about different governance strategies, they do not capture every aspect of the success of a community’s governance. What people say publicly does not always reflect their actual beliefs, and even if it did, minimizing community unhappiness is not necessarily the best objective function for community moderators. Although we conclude that excluding content that mentions specific moderators by name and removed content does not substantially impact our results (§3.4), differences in community members’ behavior and values also may still bias our results, as different community may have different norms around mod discourse, and different community members may feel more or less comfortable expressing their opinions publicly. Additional study is needed to ensure that scalable methods for measuring governance reflect the needs of all community members, not just the noisiest.
@@ 7.2. Conclusion
Good governance is critical to the functioning of online communities, yet it is difficult to know what governance practices are most effective, as it is challenging to measure the ‘success’ of community governance. In this work, we developed a method to quantify community members’ perceptions of their moderators across thousands of communities. We relate these perceptions to different aspects of governance including community size and topic (§4) as well as to different actions that moderators can take, including rule enforcement (§6.1), community engagement (§6.2), and moderator recruiting practices (§6.3-6.4). We empirically identify promising strategies for community moderators, including tailoring the strictness of rule enforcement to the community topic, and recruiting engaged mods. We make our models and anonymized datasets public to support future research.
Acknowledgements.
This research was supported by the Office of Naval Research (#N00014-21-1-2154), NSF grant IIS-1901386, NSF CAREER IIS-2142794, NSF grant CNS-2025022, and the Bill & Melinda Gates Foundation (INV-004841). This work was completed on Hyak, UW’s high performance computing cluster. This resource was funded by the UW student technology fee.
@@ 8. Ethical Considerations
We believe this work will have a positive broader impact by informing better moderation practices in online communities, as well as providing researchers with better tools to study community members’ perceptions of moderators. As we only make use of public data, we believe our work has minimal risk to participants’ privacy. As research has shown that some online community users are uncomfortable with their data being used for research, even when posted publicly (Fiesler and Proferes, 2018), we take further steps to reduce potential harms and misuse potential of our mod discourse dataset: we do not publish usernames or identifiable information, only predicted sentiment with regards to the moderators. We publish moderator timelines, including moderators’ usernames, however these usernames are already publicly listed on communities’ ‘about’ pages. Upon publication of this dataset, we will provide affordances for users to have their data removed at their request. We comply with relevant licenses for NLP models we use or modify. This study was approved by the University of Washington IRB under ID number STUDY00011457.
@@ A.3. Neutral Sentiment
Appendix B Prompts for Topic and Sentiment Classification
B.1.
The following few-shot prompt was used with GPT-4 (OpenAI et al., 2023) to classify the topic of subreddits in our analyses (§3.2), based on the name of the subreddit.
Given a subreddit, classify its topic into exactly one
of the following categories:
Hobby communities, which focus on a specific hobby,
sports related topic, or pastime.
Discussion communities, which are for discussion and
text-based content.
Media communities, which are for sharing videos,
and pictures.
News communities, which are for sharing news and
similar content.
Meme communites, which are for sharing memes or low
effort content.
Identity communities, which are for groups of a
spcecific identity or background.
/r/india: Identity
/r/bicycling,: Hobby
/r/CrappyDesign: Media
/r/me_irl: Memes
/r/worldnews: News
/r/AskReddit: Discussion
/r/nba: Hobby
/r/relationship_advice: Discussion
/r/science: News
/r/teenagers: Identity
/r/dankmemes: Memes
/r/pics: Media
/r/{subreddit}:
B.2. Sentiment Classification Step Prompt
To classify the sentiment with regards to the moderators of posts and comments discussing mods (§3.3), we used the following prompt for our LLaMA 2 model, fine tuned with QLoRA (Touvron et al., 2023; Dettmers et al., 2023).
<|im_start|>system
Given a comment from Reddit which discusses moderators (or mods),
and its parent, identify how the author of the comment feels about
the moderators of the subreddit the comment was made in.
If the author feels that the moderators are doing a good job, mark
the sentiment as positive. If the author feels the moderators are
doing a bad job, mark the sentiment as negative. If its not
possible to tell, mark the sentiment as neutral.
If the comment does not discuss moderators, but instead discusses
video game mods, or other types of modifications, mark the exclude
field as true and the sentiment as undefined.
If the comment is discussing moderators, but in a different
subreddit or different community, mark the other community
field as true.
The parent might be able to help your identification by providing
additional context.
Your answer should follow the format given in the examples.
<|im_end|>
<|im_start|>user
/r/{subreddit}
parent : {parent}
comment : {body}
<|im_end|>
Appendix C Supplementary Figures
@@
Appendix D Covariates Used in Propensity Score Modeling for IPTW
D.1. Covariates used for Mod Workload Analyses
@@
D.2. Covariates used for Strictness of Rule Enforcement Analyses